OpenStudy (anonymous):

Integrate 1 / sqrt(2x-5) dx For some reason I end up with 2(sqrt(2x-5)

myininaya (myininaya):

so i would let u=2x-5 => du=2 dx =du/2=dx

myininaya (myininaya):

\[\int\limits_{}^{}\frac{1}{\sqrt{u}} \frac{du}{2}=\frac{1}{2}\int\limits_{}^{}u^{-\frac{1}{2}} du\]

myininaya (myininaya):

\[\frac{1}{2} \cdot \frac{u^{-\frac{1}{2}+1}}{-\frac{1}{2}+1}+C\]

myininaya (myininaya):

\[\frac{1}{2} \cdot \frac{u^\frac{1}{2}}{\frac{1}{2}}+C=\sqrt{u}+C=\sqrt{2x-5}+C\]

OpenStudy (anonymous):

Thank you very much!, I missed the arrangement of du/dx at the very beggining, and it was getting me nowhere..

myininaya (myininaya):

sametimes it helps to start all the way over

myininaya (myininaya):

sometimes*